# Phase-Imaging Ion-Cyclotron-Resonance Mass Spectrometry with the Canadian Penning Trap at CARIBU

### v2

@@ Abstract
The Canadian Penning Trap mass spectrometer (CPT) has conducted precision mass measurements of neutron-rich nuclides from the CAlifornia Rare Isotope Breeder Upgrade (CARIBU) of the Argonne Tandem Linac Accelerator System (ATLAS) facility at Argonne National Laboratory for over a decade, first using Time-Of-Flight Ion-Cyclotron-Resonance (TOF-ICR) and later using Phase-Imaging Ion-Cyclotron-Resonance (PI-ICR) techniques. Here we give an overview of the CPT system as it was operated for PI-ICR measurements at CARIBU for over half a decade, along with some recently studied systematic effects.
@@
The Canadian Penning Trap (CPT) system was designed and commissioned at the Tandem Accelerator Superconducting Cyclotron (TASCC) facility at Chalk River Laboratories in Ontario, Canada [4] in the late 90s. Upon the decommissioning of TASCC, the CPT was moved to Argonne National Laboratory (ANL) [5] in 2001 and was set up in the “Triangle Room” experimental area at the Argonne Tandem Linac Accelerator System (ATLAS) facility [5]. Upon the completion of the CAlifornium Rare Isotope Breeder Upgrade (CARIBU) [6] at the ATLAS facility in 2011, the CPT was moved there [7]. Since then, it has been used to conduct precision mass measurements of neutron-rich nuclides produced at CARIBU focusing on probing the astrophysical rapid neutron capture process ( process) [8, 9] and nuclear structure features among the rare-earth nuclides [10, 11], first using Time-of-Flight Ion-Cyclotron-Resonance (TOF-ICR) [12] and then using Phase-Imaging Ion-Cyclotron-Resonance (PI-ICR) [13, 14] techniques. This fruitful period ended with its decommissioning in 2023, prior to its move to the upcoming N=126 Factory [15] at ANL. This paper discusses the CPT system as it was operated for PI-ICR in CARIBU for over half a decade, along with some new systematic studies.
@@ 2 The CARIBU facility
CARIBU is a unique facility which produces neutron-rich nuclides by collecting the spontaneous fission fragments from a 252Cf source, which are sent as low-energy beams either to the CARIBU low-energy experimental area where the CPT resided or the newly constructed Area 1, or to the Electron Beam Ion Source (EBIS) [16] for charge-breeding and acceleration before injection into ATLAS for other experiments. At the heart of CARIBU is a 0.5 Ci 252Cf source decaying with a half-life of years and having a spontaneous fission branch of 3.1028(7)% [17]. The fission fragments pass through a gold degrader foil and enter a large volume gas catcher [18] consisting of a cylindrical body and a cone. Here they are stopped and thermalized in high-purity He gas, and are extracted using a combination of gas flow, and radio-frequency (RF) and direct current (DC) fields with or . The gas catcher and associated radio-frequency quadrupole (RFQ) ion guides sit on a kV high voltage (HV) platform. The extracted ions first undergo mass-selection on the basis of their mass numbers () and by passing through the isobar separator [19] consisting of a pair of dipole magnets and operated at a resolving power of 5,000, before they are cooled and bunched using a RFQ cooler-buncher. The bunched beam is extracted at keV every 50 ms or 100 ms, and then gets further mass-resolved by a multi-reflection time-of-flight (MR-TOF) mass separator [20]. For most CPT experiments, the ions are cycled inside the MR-TOF for ms, enabling achievement of a mass resolution of 100,000. At the exit of the MR-TOF, the mass-resolved beam passes through a Bradbury-Nielson gate (BNG), where the target nuclides are precisely selected on the basis of their TOF and sent to the CPT system.
@@ 4.2 Systematic Effects
As can be seen from Eq. 1, stability of the trap’s B field is critical for accurate measurements. Spatial drifts are addressed by ensuring the ions probe the same B field in all three dimensions. Temporal drifts are handled by doing a field-calibration as close in time as possible to a target-nuclide measurement. Figure 7 shows results of -measurements conducted to gauge the stability of the B-field using 133Cs+ ions. The study was done over a period of eight days, which is the typical duration of experimental campaigns at the CPT. The measured cyclotron frequencies were found to vary statistically with a standard deviation of ppb. Additionally, several calibrant measurements with ions from CARIBU and SIS are interleaved in between target-nuclide measurements.
@@
All masses recently measured and reported were determined following these, in addition to the procedures of Ref. [14]. The uncertainty from the sine-fit systematic was incorporated in the statistical uncertainty of measurement, while that due to contaminant species was propagated from the individual components in the calculation. Furthermore, a systematic contribution of 3.2 ppb was added in quadrature to the statistical uncertainties to account for possible shifts due to magnetic field drifts, electric field instabilities, ion-ion interactions, and non-circular projection from the trap to the detector.
@@ 5 Outlook
PI-ICR implementation at the CPT at CARIBU has been a fruitful endeavour measuring masses of more than 200 neutron-rich nuclides, resulting in multiple publications and theses over six years. PI-ICR has increased the measurement sensitivity of the CPT helping achieve ppb-level precision [25, 30, 31] and record mass resolving power of 107 [32], and allowing for measurement of some of the most weakly populated 252Cf-fission fragments [8, 25]. The CPT system is currently being transported to the upcoming N=126 Factory [15] in ATLAS at ANL. Mutiple upgrades, including a revamp of the control system and introduction of a Lorentz steerer [33] just before the trap for efficient on-center injection into the trap are planned. Although this marks the end of the CPT-at-CARIBU project, it points towards an exciting future in studying neutron-rich nuclides far from stability.
@@ Acknowledgements
This work was performed with the support of US Department of Energy, Office of Nuclear Physics under Contract No. DE-AC02-06CH11357 (ANL), the Natural Sciences and Engineering Research Council of Canada under Grant No. SAPPJ-2018-00028, and the US National Science Foundation under Grant No. PHY-2011890 and PHY-2310059. This research used resources of ANL’s ATLAS facility, which is a DOE Office of Science User Facility.

### v3

@@ Abstract
The Canadian Penning Trap mass spectrometer (CPT) has conducted precision mass measurements of neutron-rich nuclides from the CAlifornia Rare Isotope Breeder Upgrade (CARIBU) of the Argonne Tandem Linac Accelerator System (ATLAS) facility at Argonne National Laboratory using the Phase-Imaging Ion-Cyclotron-Resonance (PI-ICR) technique for over half a decade. Here we discuss the CPT system, and methods to improve accuracy and precision in mass measurement using PI-ICR including some optimization techniques and recently studied systematic effects.
@@
The Canadian Penning Trap (CPT) [4, 5] is located at Argonne National Laboratory (ANL) where it has conducted precision mass measurements of neutron-rich nuclides from the CAlifornium Rare Isotope Breeder Upgrade (CARIBU) [6] at the Argonne Tandem Linac Accelerator System (ATLAS) facility, using the Phase Imaging Ion-Cyclotron-Resonance (PI-ICR) [7, 8] technique. This paper discusses the PI-ICR operations, and some new systematic studies which have enabled the CPT to achieve ppb-level precision in mass measurements.
@@ 2 The CARIBU facility
CARIBU is a unique facility which produces neutron-rich nuclides by collecting the spontaneous fission fragments from a 252Cf source, which are sent as low-energy beams either to the CARIBU low-energy experimental area where the CPT resided or the newly constructed Area 1, or to the Electron Beam Ion Source (EBIS) [9] for charge-breeding and acceleration before injection into ATLAS for other experiments. At the heart of CARIBU is a 0.5 Ci 252Cf source decaying with a half-life of years and having a spontaneous fission branch of 3.1028(7)% [10]. The fission fragments pass through a gold degrader foil and enter a large volume gas catcher [11] consisting of a cylindrical body and a cone. Here they are stopped and thermalized in high-purity He gas, and are extracted using a combination of gas flow, radio-frequency (RF) fields and direct current (DC) fields with or . The gas catcher and associated radio-frequency quadrupole (RFQ) ion guides sit on a kV high voltage (HV) platform. The extracted ions first undergo mass-selection on the basis of their mass numbers () and by passing through the isobar separator [12] consisting of a pair of dipole magnets and operated at a resolving power of 5,000, before they are cooled and bunched using a RFQ cooler-buncher. The bunched beam is extracted at keV every 50 ms or 100 ms, and then gets further mass-resolved by a multi-reflection time-of-flight (MR-TOF) mass separator [13]. For most CPT experiments, the ions are cycled inside the MR-TOF for ms, enabling achievement of a mass resolution of 100,000. At the exit of the MR-TOF, the mass-resolved beam passes through a Bradbury-Nielson gate (BNG), where the target nuclides are precisely selected on the basis of their TOF and sent to the CPT system.
@@ 4.2 Systematic Effects
As can be seen from Eq. 1, stability of the trap’s B field is critical for accurate measurements. Spatial drifts are addressed by ensuring the ions probe the same B field in all three dimensions. Temporal drifts are handled by doing a field-calibration as close in time as possible to a target-nuclide measurement. Figure 7 shows results of -measurements conducted to gauge the stability of the B-field using 133Cs+ ions. The study was done over a period of eight days, which is the typical duration of experimental campaigns at the CPT. The measured cyclotron frequencies were fitted to a linear model with a slope of 0.0(6) ppbday. During mass measurements, several calibrant measurements with ions from CARIBU and SIS are interleaved in between target-nuclide measurements. Additionally, possible mass shifts from magnetic field drifts are calculated and added in quadrature to the measured uncertainties.
@@
All masses recently measured and reported were determined following these, in addition to the procedures of Ref. [8]. The uncertainty from the sine-fit systematic was incorporated in the statistical uncertainty of measurement, while that due to contaminant species was propagated from the individual components in the calculation. Furthermore, a systematic contribution of 3.0 ppb was added in quadrature to the statistical uncertainties to account for possible shifts due to electric field instabilities, ion-ion interactions, and non-circular projection from the trap to the detector, in addition to the magnetic field shift.
@@ 5 Outlook
PI-ICR implementation at the CPT at CARIBU has been a fruitful endeavour measuring masses of more than 200 neutron-rich nuclides over six years, resulting in multiple publications and theses. PI-ICR has increased the measurement sensitivity of the CPT helping achieve ppb-level precision [18, 23, 24, 25] and record mass resolving power of 107 [26], and allowing for measurement of some of the most weakly populated 252Cf-fission fragments [27, 18]. The CPT system is currently being transported to the upcoming N=126 Factory [28] in ATLAS at ANL. While most of the system will be kept unchanged, a few upgrades, including a revamp of the control system and introduction of a Lorentz steerer [29] just before the trap for efficient on-center injection into the trap are planned. Although this marks the end of the CPT-at-CARIBU project, it points towards an exciting future in studying neutron-rich nuclides far from stability.
@@ Acknowledgements
This work was performed with the support of US Department of Energy, Office of Nuclear Physics under Contract No. DE-AC02-06CH11357 (ANL), the Natural Sciences and Engineering Research Council of Canada under Grant No. SAPPJ-2018-00028, and the US National Science Foundation under Grant No. PHY-2310059. This research used resources of ANL’s ATLAS facility, which is a DOE Office of Science User Facility.